# Less is More: Local Intrinsic Dimensions of Contextual Language Models

### v1

@@ Abstract
Understanding the internal mechanisms of large language models (LLMs) remains a challenging and complex endeavor. Even fundamental questions, such as how fine-tuning affects model behavior, often require extensive empirical evaluation. In this paper, we introduce a novel perspective based on the geometric properties of contextual latent embeddings to study the effects of training and fine-tuning. To that end, we measure the local dimensions of a contextual language model’s latent space and analyze their shifts during training and fine-tuning. We show that the local dimensions provide insights into the model’s training dynamics and generalization ability. Specifically, the mean of the local dimensions predicts when the model’s training capabilities are exhausted, as exemplified in a dialogue state tracking task, overfitting, as demonstrated in an emotion recognition task, and grokking, as illustrated with an arithmetic task. Furthermore, our experiments suggest a practical heuristic: reductions in the mean local dimension tend to accompany and predict subsequent performance gains. Through this exploration, we aim to provide practitioners with a deeper understanding of the implications of fine-tuning on embedding spaces, facilitating informed decisions when configuring models for specific applications. The results of this work contribute to the ongoing discourse on the interpretability, adaptability, and generalizability of LLMs by bridging the gap between intrinsic model mechanisms and geometric properties in the respective embeddings.
@@ 1 Introduction
Contributions. We present a unified framework for analyzing LLM training dynamics through dimensionality. Based on this framework, we (i) demonstrate that fine-tuning reshapes the local intrinsic dimension in a dataset-specific manner, allowing us to infer overlap between training data used for fine-tuning with validation (or test) data from geometric shifts alone, without requiring labels; (ii) show that dimensionality anticipates the onset of grokking in synthetic arithmetic tasks, identifying generalization beyond training data from the dimensionality of the model’s embeddings of the training data alone; (iii) demonstrate that local intrinsic dimensions can predict training convergence, as evidenced in a sequence-tagging-based dialogue state tracking task where stabilizing dimension correlates with model stabilization; and (iv) show that local intrinsic dimensions detect overfitting in a sequence classification setting, where an initial drop followed by a rise in dimension reflects the model’s tradeoff between generalization and memorization. In total, our results suggest that local intrinsic dimensions serve as a valuable unsupervised signal for practitioners seeking to interpret and monitor LLM behavior. Beyond offering empirical insights, our work thus highlights the potential of geometric descriptors to complement traditional evaluation methods and potentially serve to inform future model design.
@@ 2 Related Work
In a first qualitative analysis of the internal representations of transformer models, Ethayarajh, (2019) and Cai et al., (2021) identify clusters and low-dimensional manifold structures in contextual embedding spaces. Recent work has started studying relationships between geometric properties of LLMs and corresponding semantics. In Tulchinskii et al., (2024), the authors discover that the average intrinsic global dimension of artificially generated texts is lower than that of human-written texts. In that work, the points in the latent space are taken from a single text paragraph, so their work applies to single data input sequences, not to entire datasets like our approach. Valeriani et al., (2024) investigate how the global intrinsic dimension estimated is altered as data is passed through an LLM. Aghajanyan et al., (2021) define a notion of intrinsic dimension based on restricting the model’s parameter space and its effect on the objective function, and show that larger LLMs tend to have smaller intrinsic dimensionality. That work studies the dimensionality of the model’s parameter space, and not the latent space created from specific datasets. In Viswanathan et al., (2025), the authors analyze token-level intrinsic dimensions, linking token geometry to model next-token-prediction loss. The difference between their method and ours is that they always consider the contextual token embeddings of a single, sufficiently long prompt as making up the embedding space. In contrast, we sub-sample from an entire dataset split.
@@
The topological descriptors in Kushnareva et al., (2021); Perez and Reinauer, (2022); Tulchinskii et al., (2024) correspond to entire sequences of input data, while our proposed estimates are defined on the token level. In Durrani et al., (2022), pre-trained and fine-tuned models are compared using hierarchical clustering and alignment functions. The results indicate that the latent space in the higher layers adapts to task-specific concepts, while the lower layers preserve the general concepts learned in the pre-trained model. Ed-dib et al., (2024) introduces a method for fine-tuning large language models by dynamically adjusting LoRA ranks based on the intrinsic dimensionality of hidden state representations. Here, the intrinsic dimension is measured as the rank of an information matrix, and not from the data space as in our work. In Chang et al., (2022), the representational geometry of multilingual language models is explored, focusing on how they balance encoding language-specific and language-agnostic features within a shared multilingual space. Ruppik et al., (2024) utilize topological features derived from neighborhoods in a contextual embedding space to improve performance in a sequence tagging task. The definition of their features requires a fixed ambient data corpus, whereas our methods give an intrinsic measure of the dataset under consideration; thus, our measures are comparable between different model checkpoints of the feature generation model. In contrast to mechanistic interpretability (see, for example, the survey (Ferrando et al.,, 2024)), in which one tries to explain how a language model’s internal representations fit together into higher-level abstractions, our approach is more atomic. We study the smallest non-divisible representation in the model (the contextual embedding of single tokens) and the space created by collections of these.
@@ 3.2 Latent Space Modeling
with $I_{m}$ denoting token indices in the $m$-th sequence. Distances between points are measured in the ambient Euclidean space and are assumed to be Euclidean unless stated otherwise. In practice, $\mathbb{T}_{i}$ can contain millions of vectors, making subsequent neighborhood computation on the full dataset infeasible.
In order to obtain a representative subset of tokens $\mathbb{T}$, we take a two-step sampling approach: We first sample $M$ sequences from $\mathcal{D}$, and after de-duplication, we sample $N$ vectors from the resulting token embedding space. Finally, we compute neighborhoods $\mathcal{N}_{L}(t_{j};\mathbb{T})$ for each token using a locality parameter $L$. We assume the dataset is drawn from an underlying text-generating distribution, and our sampling approximates its geometric structure. We confirm this assumption in our sensitivity analysis in Appendix A.
@@
1 Take a random sub-sample $\mathcal{D}_{\sim M}$ of size $M$ of the dataset sequences ;
@@ 3.4 Local Dimension Estimates
We turn this into a local measure as follows: Let $X$ be a point cloud and $v\in X$. For $L\in\mathbb{N}_{>0}$, we define the $L$-local TwoNN estimate of $X$ at $v$ as $\operatorname{TwoNN}(\mathcal{N}_{L}(v;X))$, i.e. as the TwoNN estimate of the local neighborhood $\mathcal{N}_{L}(v;X)$ of $v$ of size $L$ in $X$. In this way, the parameter $L$ controls the locality scale of the dimension estimation. In the context of our model $\mathcal{M}$ and corpus $\mathcal{D}$, we take $X=\mathbb{T}$ and compute a vector of positive real-valued numbers $\operatorname{TwoNN}(\mathcal{N}_{L}(v;X))_{v\in\mathbb{T}}\in\mathbb{R}^{N}_{\geq 0}$, one for each token $v\in\mathbb{T}$. We subsequently aggregate this to a mean estimate per corpus/model setup. While sensitive to hyperparameters such as $L$ and sample sizes, we show in Appendix A that estimates remain stable under reasonable settings, enabling consistent comparisons across datasets and models.
@@ 4 Experiments
We now apply our method to analyze how embedding spaces evolve during LLM-related learning tasks. Unless stated otherwise, all results pertain to the model’s final hidden layer, i.e., $\mathbb{T}_{-1}$ as defined in Section 3. We refer the reader to Section C.2 for results on layers other than the last. Subsequently, we will focus on four central questions:
@@
    How can local dimension estimates predict grokking?
@@ 4.1 Fine-Tuning Induces Dataset-Specific Shifts in Heterogeneous Local Dimensions
To understand how fine-tuning alters model representations, we investigate the distribution of local intrinsic dimensions across token embeddings. These dimensions, estimated using the method introduced in Section 3, reveal nuanced changes in the geometry of the embedding space.
@@
We evaluate models on datasets with varying domain and familiarity to the base and fine-tuned models: MultiWOZ2.1 (Eric et al.,, 2020): Human-human multi-domain dialogues, Schema-Guided Dialogue (SGD) (Rastogi et al.,, 2020): Human–virtual assistant dialogues, Reddit: 2022 Reddit comments mentioning Tesla, Inc., and Wikipedia: The Hugging Face wikitext-103-v1 corpus. As can be seen from the pre-training datasets of the publications that introduced the respective models (Liu et al.,, 2019; Radford et al.,, 2019), this selection allows us to compare: (i) distributions not seen during pre-training or fine-tuning (e.g., Reddit), (ii) distributions seen during pre-training (Wikipedia), and (iii) distributions used for fine-tuning (e.g., MultiWOZ, SGD). We split Reddit and Wikipedia into training (80%), validation (10%), and test (10%) subsets. Chronologically-ordered datasets (e.g., Reddit) are shuffled before splitting to avoid temporal bias. Wikipedia was pre-processed by removing headings, empty lines, and stripping leading and trailing whitespace. Further dataset statistics are available in Table 1.
Fine-tuning of the RoBERTa-base models (Liu et al.,, 2019) is performed using masked language modeling with a masking probability of $0.15$. Each model is trained for 5 epochs on 10 0001000010\,00010 000 training examples using a batch size of 8, a learning rate peaking at $5\cdot 10^{-5}$ with 500 warmup steps, and linear decay thereafter. Weight decay of $0.01$ is applied throughout. For evaluation, we embed a subset of the validation split of each dataset using both the base and fine-tuned model, ensuring the embeddings are out-of-sample relative to the fine-tuning set. Results for fine-tuning an autoregressive model can be found in Section C.1.
(a) TwoNN estimates on MultiWOZ validation.
(b) TwoNN estimates on Wikipedia validation.
(c) TwoNN estimates on Reddit validation.
Figure 1: Comparison of the intrinsic dimension (ID) across three data modalities for both models. The distribution of the local estimates over tokens is shown in the violin plot, together with their means and quartiles. The ID of embeddings originating from the fine-tuning distribution (MultiWOZ) differs markedly between models, whereas the IDs for the out-of-distribution corpora (Wikipedia, Reddit) are almost indistinguishable.
@@ Results
[…]
Our central observation is that fine-tuning systematically lowers the local intrinsic dimension only on the dataset used during the fine-tuning process. As shown in Figure 1a, the TwoNN estimates for the MultiWOZ-validation embeddings are markedly lower for the fine-tuned RoBERTa model than for the base model. By contrast, embeddings of unrelated datasets—such as Wikipedia (Figure 1b) or Reddit comments (Figure 1c)—show no or only minimal changes in their local dimension distributions. Quantitatively, this observation can be validated by the standardized mean difference between the respective cohorts, with a value of $1.19$ for MultiWOZ, and $0.08$ (resp. $0.1$) for Reddit (resp. Wikipedia).
We find this shift to be stable across training, validation, and test splits of the same dataset, suggesting that it reflects robust structural changes in the latent space rather than artifacts (see Appendix A for a thorough sensitivity analysis showing the method’s robustness in our setting). This behavior highlights strong dataset-specificity in how local geometry adapts under fine-tuning. The effect is absent when evaluating on out-of-distribution data, supporting the view that local dimension reductions correspond to improved fit and specialization on the fine-tuned task, while leaving unrelated regions of the embedding space unchanged.
4.2 Local Dimensions Predict Grokking
@@
Figure 2: Training a model on addition mod $p=197$ with different training data fraction selected from $\{0.1;0.15;0.2;0.25;0.3;0.4;0.5\}$. The plots show the development for 60 0006000060\,00060 000 batches, with mean and 95% confidence interval over $5$ training seeds per configuration (plots per seed are in Section C.3). The mean local estimates are computed on the training split for the parameters $N=3000$; $L=64$. Dashed lines highlight the runs where grokking did not occur.
@@ Results
Performance measures versus local estimates are shown in Figure 2 for the first 60 0006000060\,00060 000 training steps averaged over seeds (we present plots for individual seeds in Figure 15 in the appendix). The training accuracy quickly reaches an almost perfect score in all cases, with the validation accuracy lagging behind. In the time frame under consideration, the model can generalize and reach almost perfect validation accuracy only in those runs for which the training data portion exceeds 20%. Based on the loss and accuracy of the training data alone, one would not be able to predict which configurations can break out and generalize to the validation data.
A typical pattern for the mean local dimension computed on the training data in all these runs is that it increases in the first few thousand global steps (see Figure 2(c)). But subsequently, the training mean local dimension starts dropping significantly for those runs that exhibit grokking. Observe that the timing of this drop coincides with the start of the increasing validation accuracy in Figure 2(b). We can conclude that in this setting, a drop in the mean local dimension on the training set strongly indicates a successful generalization to unseen validation data.
For those two runs in which the validation accuracy does not increase beyond the fraction of training data during the selected training time (with 10% and 15% of training data, highlighted via dashed lines), the local dimension increases and then stays mostly flat. This behavior might hint that the model only tries to learn the training examples by heart and fails to generalize to unseen data.

### v2

@@ Abstract
Understanding the internal mechanisms of large language models (LLMs) remains a challenging and complex endeavor. Even fundamental questions, such as how fine-tuning affects model behavior, often require extensive empirical evaluation. In this paper, we introduce a novel perspective based on the geometric properties of contextual latent embeddings to study the effects of training and fine-tuning. To that end, we measure the local dimensions of a contextual language model’s latent space and analyze their shifts during training and fine-tuning. We show that the local dimensions provide insights into the model’s training dynamics and generalization ability. Specifically, the mean of the local dimensions predicts when the model’s training capabilities are exhausted, as exemplified in a dialogue state tracking task, overfitting, as demonstrated in an emotion recognition task, and grokking, as illustrated with an arithmetic task. Furthermore, our experiments suggest a practical heuristic: reductions in the mean local dimension tend to accompany and predict subsequent performance gains. Through this exploration, we aim to provide practitioners with a deeper understanding of the implications of fine-tuning on embedding spaces, facilitating informed decisions when configuring models for specific applications. The results of this work contribute to the ongoing discourse on the interpretability, adaptability, and generalizability of LLMs by bridging the gap between intrinsic model mechanisms and geometric properties in embeddings.
@@ 1 Introduction
Contributions We present a unified framework for analyzing LLM training dynamics through dimensionality. Based on this framework, we (i) demonstrate that fine-tuning reshapes the local intrinsic dimension in a dataset-specific manner, allowing us to infer overlap between training data used for fine-tuning with validation (or test) data from geometric shifts alone, without requiring labels; (ii) show that dimensionality anticipates the onset of grokking in synthetic arithmetic tasks, identifying generalization beyond training data from the dimensionality of the model’s embeddings of the training data alone; (iii) demonstrate that local intrinsic dimensions can indicate training convergence, as evidenced in a sequence-tagging-based dialogue state tracking task where stabilizing dimension correlates with model stabilization; and (iv) show that local intrinsic dimensions detect overfitting in a sequence classification setting, where an initial drop followed by a rise in dimension reflects the model’s tradeoff between generalization and memorization. In total, our results suggest that local intrinsic dimensions serve as a valuable unsupervised signal for practitioners seeking to interpret and monitor LLM behavior. Beyond offering empirical insights, our work thus highlights the potential of geometric descriptors to complement traditional evaluation methods and potentially serve to inform future model design. [^1] Our code is available at https://github.com/aidos-lab/Topo_LLM_public and https://github.com/aidos-lab/grokking-via-lid.
@@ 2 Related Work
In a first qualitative analysis of the internal representations of transformer models, Ethayarajh, (2019) and Cai et al., (2021) identify clusters and low-dimensional manifold structures in contextual embedding spaces. Recent work has started studying relationships between geometric properties of LLMs and corresponding semantics. In Tulchinskii et al., (2023), the authors discover that the average intrinsic global dimension of artificially generated texts is lower than that of human-written texts. In that work, the points in the latent space are taken from a single text paragraph, so their work applies to single data input sequences, not to entire datasets like our approach. Valeriani et al., (2023) investigate how the global intrinsic dimension estimated is altered as data is passed through an LLM. Aghajanyan et al., (2021) define a notion of intrinsic dimension based on restricting the model’s parameter space and its effect on the objective function, and show that larger LLMs tend to have smaller intrinsic dimensionality. That work studies the dimensionality of the model’s parameter space, and not the latent space created from specific datasets. In Viswanathan et al., (2025), the authors analyze token-level intrinsic dimensions, linking token geometry to model next-token-prediction loss. The difference between their method and ours is that they always consider the contextual token embeddings of a single, sufficiently long prompt as making up the embedding space. In contrast, we sub-sample from an entire dataset split. Lee et al., 2025a identify structural similarities between the token embedding spaces of different language models, particularly reporting correlations in local intrinsic dimension measures. However, their analysis is limited to the non-contextual token embedding and unembedding layers, without considering contextual representations.
@@
The topological descriptors in Kushnareva et al., (2021); Perez and Reinauer, (2022); Tulchinskii et al., (2023); Lee et al., 2025b correspond to entire sequences of input data, while our proposed estimates are defined on the token level. In Durrani et al., (2022), pre-trained and fine-tuned models are compared using hierarchical clustering and alignment functions. The results indicate that the latent space in the higher layers adapts to task-specific concepts, while the lower layers preserve the general concepts learned in the pre-trained model. Ed-dib et al., (2024) introduces a method for fine-tuning large language models by dynamically adjusting LoRA ranks based on the intrinsic dimensionality of hidden state representations. Here, the intrinsic dimension is measured as the rank of an information matrix, and not from the data space as in our work. In Chang et al., (2022), the representational geometry of multilingual language models is explored, focusing on how they balance encoding language-specific and language-agnostic features within a shared multilingual space. Ruppik et al., (2024) utilize topological features derived from neighborhoods in a contextual embedding space to improve performance in a sequence tagging task. The definition of their features requires a fixed ambient data corpus, whereas our methods give an intrinsic measure of the dataset under consideration; thus, our measures are comparable between different model checkpoints of the feature generation model. Both works are situated in the nascent field of topological deep learning (Papamarkou et al.,, 2024), employing concepts from geometry and topology in an observational manner to study machine learning models. In contrast to mechanistic interpretability (see, for example, the survey by Ferrando et al., 2024), in which one tries to explain how a language model’s internal representations fit together into higher-level abstractions, our approach is more atomic. We study the smallest non-divisible representation in the model (the contextual embedding of single tokens) and the space created by collections of these.
@@ 3.2 Latent Space Modeling
with $I_{m}$ denoting token indices in the $m$-th sequence. Distances between points are measured in the ambient Euclidean space. In practice, $\mathbb{T}_{i}$ can contain millions of vectors, making subsequent neighborhood computation on the full dataset infeasible.
To obtain a representative subset of tokens $\mathbb{T}$, we take a two-step sampling approach: We first sample $M$ sequences from $\mathcal{D}$, and after de-duplication, we sample $N$ vectors from the resulting token embedding space. Finally, we compute neighborhoods $\mathcal{N}_{L}(t_{j};\mathbb{T})$ for each token using a locality parameter $L$. We assume the dataset is drawn from an underlying text-generating distribution, and our sampling approximates its geometric structure. We confirm this assumption in our sensitivity analysis in Appendix˜A.
@@
1
21exTake a random sub-sample $\mathcal{D}_{\sim M}$ of size $M$ of the dataset sequences ;
@@ 3.4 Local Dimension Estimates
We turn this into a local measure as follows: Let $X$ be a point cloud and $v\in X$. For $L\in\mathbb{N}_{>0}$, we define the $L$-local TwoNN estimate of $X$ at $v$ as $\operatorname{TwoNN}(\mathcal{N}_{L}(v;X))$, i.e., as the TwoNN estimate of the local neighborhood $\mathcal{N}_{L}(v;X)$ of $v$ of size $L$ in $X$. In this way, the parameter $L$ controls the locality scale of the dimension estimation. In the context of our model $\mathcal{M}$ and corpus $\mathcal{D}$, we take $X=\mathbb{T}$ and compute a vector of positive real-valued numbers $\operatorname{TwoNN}(\mathcal{N}_{L}(v;X))_{v\in\mathbb{T}}\in\mathbb{R}^{N}_{\geq 0}$, one for each token $v\in\mathbb{T}$. We subsequently aggregate this to a mean estimate per corpus/model setup. While sensitive to hyperparameters such as $L$ and sample sizes, we show in Appendix˜A that estimates remain stable under reasonable settings, enabling consistent comparisons across datasets and models.
@@ 4 Experiments
We now apply our method to analyze how embedding spaces evolve during LLM-related learning tasks. Unless stated otherwise, all results pertain to the model’s final hidden layer, i.e., $\mathbb{T}_{-1}$ as defined in Section˜3. We refer the reader to Section˜C.2 for results on layers other than the last. Subsequently, we will focus on four central questions:
@@
    How can local dimension estimates detect grokking?
@@ 4.1 Fine-Tuning Induces Dataset-Specific Shifts in Heterogeneous Local Dimensions
To understand how fine-tuning alters model representations, we investigate the distribution of local intrinsic dimensions across token embeddings. These dimensions, estimated using the method introduced in Section˜3, reveal nuanced changes in the geometry of the embedding space.
(a) TwoNN estimates on MultiWOZ validation
(b) TwoNN estimates on Wikipedia validation
(c) TwoNN estimates on Reddit validation
Figure 1: Comparison of local intrinsic dimensions (LIDs) across three data modalities. The distribution of the local estimates over tokens is shown in the violin plot, together with their means and quartiles. The LID of embeddings originating from the fine‑tuning distribution (MultiWOZ) differs markedly between models, whereas the LIDs for the out‑of‑distribution corpora (Wikipedia, Reddit) are almost indistinguishable.
@@
We evaluate models on datasets with varying domains and familiarity to the base and fine-tuned models: MultiWOZ2.1 (Eric et al.,, 2020): Human-human multi-domain dialogues; Schema-Guided Dialogue SGD (Rastogi et al.,, 2020): Human–virtual assistant dialogues; Reddit: Reddit comments from the year 2022 mentioning Tesla, Inc.; ICLR 2024 Submissions: Titles and abstracts of ICLR 2024 papers collected by us; and Wikipedia: The Hugging Face wikitext-103-v1 corpus.
As can be seen from the selection of pre-training datasets of the respective models (Liu et al.,, 2019; Radford et al.,, 2019), this choice allows us to compare: (i) distributions not seen during pre-training or fine-tuning (e.g., Reddit, ICLR 2024, which were released after the pre-training of the model concluded), (ii) distributions seen during pre-training (Wikipedia), and (iii) distributions used for fine-tuning (e.g., MultiWOZ, SGD). This contrast allows us to probe how local intrinsic dimensions behave in both seen and unseen data regimes. We split Reddit and Wikipedia into training (80%), validation (10%), and test (10%) subsets. Chronologically-ordered datasets (e.g., Reddit) are shuffled before splitting to avoid temporal bias. Wikipedia was pre-processed by removing headings, empty lines, and stripping leading and trailing whitespace. Further dataset statistics are available in Table˜1.
Fine-tuning of the RoBERTa-base models (Liu et al.,, 2019) is performed using masked language modeling with a masking probability of $0.15$. Each model is trained for 5 epochs on 10 00010\,000 training examples using a batch size of 8, a learning rate peaking at $5\cdot 10^{-5}$ with 500 warmup steps, and linear decay thereafter. Weight decay of $0.01$ is applied throughout. For evaluation, we embed a subset of the validation split of each dataset using both the base and fine-tuned model, ensuring the embeddings are out-of-sample relative to the fine-tuning set.
Our discussion in this section focuses on the masked language model RoBERTa, whose architectures remain widely used in practice where encoder-based models are needed (Warner et al.,, 2025) and autoregressive alternatives are less suitable, such as in information retrieval. Results for additional fine-tuning of the autoregressive models GPT-2-medium, Phi-3.5-mini-instruct, and Llama-3.1-8B can be found in Section˜C.1, where similar observations hold.
@@ Results
We find that local intrinsic dimensions vary significantly across tokens, as illustrated by the wide spread of the TwoNN distributions in Figure˜1. Here, we choose the parameters M=7000M=\$7000\$, N=60  [^0] N=\$60\,000\$, L=128L=\$128\$. This heterogeneity is consistent across different models and datasets. It aligns with prior observations that numerical data is often not confined to a single manifold of uniform dimensionality, but is instead composed of multiple regions with varying local geometry (Brown et al.,, 2023). These findings reinforce our decision to favor local over global estimates of embedding dimensionality.
Our central observation is that fine-tuning systematically lowers the local intrinsic dimension only on the dataset used during the fine-tuning process. As shown in Figure˜1a, the TwoNN estimates for the MultiWOZ-validation embeddings are markedly lower for the fine-tuned RoBERTa model than for the base model. By contrast, embeddings of unrelated datasets—such as Wikipedia (Figure˜1b) or Reddit comments (Figure˜1c)—show no or only minimal changes in their local dimension distributions. Quantitatively, this observation can be validated by the standardized mean difference between the respective cohorts, with a value of $1.19$ for MultiWOZ, and $0.08$ (resp. $0.1$) for Reddit (resp. Wikipedia).
We find this shift to be stable across training, validation, and test splits of the same dataset, suggesting that it reflects robust structural changes in the latent space rather than artifacts (see Appendix˜A for a thorough sensitivity analysis showing the method’s robustness to the choice of sampling parameters). This behavior highlights strong dataset-specificity in how local geometry adapts under fine-tuning. The effect is absent when evaluating on out-of-distribution data, supporting the view that local dimension reductions correspond to improved fit and specialization on the fine-tuned task, while leaving unrelated regions of the embedding space unchanged.
4.2 Local Dimensions Detect Grokking
@@
Figure 2: Training a model on addition mod $p=197$ with different training data fraction selected from $\{0.1;0.15;0.2;0.25;0.3;0.4;0.5\}$. The plots show the development for 60 00060\,000 batches, with mean and 95% confidence interval over 55 training seeds per configuration (plots per seed are in Section˜C.3). The mean local estimates are computed on the training split for the parameters $N=3000$; $L=64$. Dashed lines highlight the runs where grokking did not occur.
@@ Results
Performance measures versus local estimates are shown in Figure˜2 for the first 60 00060\,000 training steps averaged over seeds (we present plots for individual seeds in Figure˜16 in the appendix). The training accuracy quickly reaches an almost perfect score in all cases, with the validation accuracy lagging behind. In the time frame under consideration, the model can generalize and reach almost perfect validation accuracy only in those runs for which the training data portion exceeds 20%. Based on the loss and accuracy of the training data alone, one would not be able to predict which configurations can break out and generalize to the validation data.
A typical pattern for the mean local dimension computed on the training data in all these runs is that it increases in the first few thousand global steps (see Figure˜2(c)). But subsequently, the training mean local dimension starts dropping significantly for those runs that exhibit grokking. Observe that the timing of this drop coincides with the start of the increasing validation accuracy in Figure˜2(b). We can conclude that in this setting, a drop in the mean local dimension on the training set strongly indicates a successful generalization to unseen validation data.
For those runs in which the validation accuracy does not increase beyond the fraction of training data during the selected training time (with 10% and 15% of training data, highlighted via dashed lines), the local dimension increases and then stays mostly flat. Quantitatively, the Spearman rank correlation between the training/validation accuracy and mean local dimension measured over the training split is positive for those training fractions where the model does not grok ($0.880/0.076$ and $0.922/0.257$ for training fractions 10% and 15%). This behavior suggests that the model only tries to learn the training examples by heart and fails to generalize to unseen data. For all runs grouped by training fraction $\geq 20$%, which all exhibit grokking, the Spearman rank correlation between both training/validation accuracy and training mean local dimension is negative.